OpenStudy (anonymous):

Joelle earns her regular pay of $7.50 per hour for up to 40 hours of work in a week. For each hour over 40 hours in a week, she is paid 1 and 1/2 times her regular pay. how much does joelle earn for a week in which she works 42 hours?

OpenStudy (anonymous):

is the answer $472.50?

OpenStudy (anonymous):

$7.50 an hour = if she works up to 40 hours $7.50 x 1.5 = if she works more than 40 hours.

OpenStudy (anonymous):

no it is not.

OpenStudy (anonymous):

please tell me your working out, how you achieved that answer.

OpenStudy (anonymous):

I multiplied 7.50 by 1.5=11.25 then i multiplied 11.25 by 42 which equals 472.50

OpenStudy (anonymous):

now that is wrong because that only applies for when she works for more than 40 hours.

OpenStudy (anonymous):

she did. she worked 42 hours

OpenStudy (anonymous):

nononono.

OpenStudy (anonymous):

ok. explain

OpenStudy (anonymous):

"More than" 40 hours key word "More than"

jimthompson5910 (jim_thompson5910):

ashleyb that answer would be correct if she got paid 1 and a half times her regular pay of $7.50 for the entire 42 hours but that overtime bonus only applies for the additional 2 hrs (not the full 42 hrs)

OpenStudy (anonymous):

$7.50 an hour = if she works up to 40 hours $7.50 x 1.5 = if she works more than 40 hours. How many hours did she work for more than 40 hours? so 40 + 2 = 42 hours that she worked in total. do you see she has worked 2 hours over 40 hours?

OpenStudy (stamp):

\[r_{regular}=750\ cents\]\[r_{overtime}=150\%(r_{regular})\]\[r_{overtime}=\frac{150}{100}750=15*75=750+5(75)=\]\[750+375=1125\frac{cents}{hour}=r_{overtime}\] pay = regular + overtime \[pay_{cents}=750(40)+1125(2)\]

jimthompson5910 (jim_thompson5910):

you would have to do it in pieces calculate her total pay for working 40 hrs at $7.50 an hour then add on the additional bonus pay of working 2 hrs at 1.5*7.50 = 11.25 dollars an hour

OpenStudy (anonymous):

yes @jim_thompson5910 says it well, she has worked additionally two hours on top of the 40 base hours.

OpenStudy (anonymous):

jayds: yes

OpenStudy (anonymous):

311.25?

OpenStudy (anonymous):

yes u are right.

OpenStudy (anonymous):

wait.

OpenStudy (anonymous):

i think u may have mistyped it.

jimthompson5910 (jim_thompson5910):

I'm getting 322.5, so you're 11.25 off somehow

OpenStudy (anonymous):

in the calculator.

jimthompson5910 (jim_thompson5910):

you must have only calculated 1 additional hour of overtime instead of 2

OpenStudy (stamp):

300 + 22.50 = 322.50 $

OpenStudy (anonymous):

First 40 hours = base pay of $7.50 per hour additional 2 hours = base pay x 1.5 = 7.50 x 1.5 = 11.25 per hour

OpenStudy (anonymous):

stamp: how did u get 322.50

jimthompson5910 (jim_thompson5910):

Regular: 40 hours at $7.50 an hour = $300 in total pay Overtime: 2 hours at $11.25 an hour = $22.50 in total pay Overall, she made 300+22.50 = 322.50 dollars working 42 hours

OpenStudy (anonymous):

OHHHHH!!!! i understand!! thankyou!

jimthompson5910 (jim_thompson5910):

you're welcome